find the measure of the supplement of angle p, where the measure of angle p= (6y+40) degrees
Answer by nyc_function(2741)

The supplement of p = (180 - y).

180 - y = 6y + 40

180 - 40 = 6y + y

140 = 7y

140/7 = y

20 = y

Then 180 - 20 = 160 degrees.